A population consists of the five measurements 2, 6, 8, 0, and 1.

1. What is the mean and the standard deviation of the population? 



1
Expert's answer
2022-04-19T04:29:10-0400

1. Mean



"\mu=\dfrac{2+6+8+0+1}{5}=3.4"

Variance



"\sigma^2=\dfrac{1}{5}\big((2-3.4)^2+(6-2.4)^2+(8-3.4)^2""(0-3.4)^2+(1-3.4)^2\big)=9.44"


Standard deviation



"\sigma=\sqrt{\sigma^2}=\sqrt{9.44}\approx3.0725"